Equation = Sin (t + 12) = sin (90)
Question; solve t make t the subject
1
Expert's answer
2020-04-26T09:41:42-0400

We have an equation


"\text{sin} (t + 12) = \text{sin} (90)."


The values of two sines can be equal only when their arguments (the part inside the parentheses) are equal, therefore, we can put that


"t+12=90,\\
t=90-12=78."

The answer is "t=78".